# Are there identifiable structural parts in the sentence embedding whole?

### v1

@@ Abstract
Sentence embeddings from transformer models encode in a fixed length vector much linguistic information. We explore the hypothesis that these embeddings consist of overlapping layers of information that can be separated, and on which specific types of information – such as information about chunks and their structural and semantic properties – can be detected. We show that this is the case using a dataset consisting of sentences with known chunk structure, and two linguistic intelligence datasets, solving which relies on detecting chunks and their grammatical number, and respectively, their semantic roles, and through analyses of the performance on the tasks and of the internal representations built during learning. [^1] The work was done while the authors were at the University of Geneva
@@ 4.1.1 Experimental set-up
The architecture of the sentence-level VAE is similar to a previously proposed system Nastase and Merlo (2023). The encoder consists of a CNN layer with a 15x15 kernel, which is applied to a 32x24-shaped sentence embedding, [^6] Nastase and Merlo (2023) show that task-relevant information is more easily accessible in transformer-based sentence embeddings reshaped as two-dimensional arrays. followed by a linear layer that compresses the output of the CNN into a latent layer of size 5. The decoder is a mirror-image of the encoder, and unpacks a sampled latent vector into a 32x24 sentence representation.
@@ 4.2.2 Analysis
Second, from the results in terms of F1, we note that the two-level process generalizes better from simpler data – learning on type I and type II leads to better results on all test data, with the highest improvement when tested on type III data, which has the highest lexical variation. Furthermore, the two-level models learned when training on the lexically simpler data perform better when tested on the type III data than the models learned on type III data itself. This result not only indicates that structure information is more easily detectable when lexical variation is less of a factor, but more importantly, that chunk information is separable from other types of information in the sentence embedding, as the patterns detecting it can be applied successfully for data with additional (lexical) variation. [^7] It might appear surprising that the two-level approach leads to lower performance on type III data, particularly when lexical variation had not been an issue for the sentence representation analysis (Section 4.1). The difference comes from the way the instances were formed, on the fly, for the two-level process: the positive sentence to be reconstructed is the same as the input, instead of being a sentence that has the same structure, but different lexical material. This is because all sentences in the sequence have different structures. We think this weakens the (indirect) supervision signal – as the correct answer is distinct from the other options. This is not the case for type I and II data, where, because of the very similar lexical material, the distinction between the correct and incorrect answers reduce to the structure. We plan to confirm this in future work using a pre-trained sentence-level VAE.

### v2

@@ Abstract
Sentence embeddings from transformer models encode in a fixed length vector much linguistic information. We explore the hypothesis that these embeddings consist of overlapping layers of information that can be separated, and on which specific types of information – such as information about chunks and their structural and semantic properties – can be detected. We show that this is the case using a dataset consisting of sentences with known chunk structure, and two linguistic intelligence datasets, solving which relies on detecting chunks and their grammatical number, and respectively, their semantic roles, and through analyses of the performance on the tasks and of the internal representations built during learning.  [^1] The work was done while the authors were at the University of Geneva.
@@ 4.1.1 Experimental set-up
The architecture of the sentence-level VAE is similar to a previously proposed system Nastase and Merlo (2023). The encoder consists of a CNN layer with a 15x15 kernel, which is applied to a 32x24-shaped sentence embedding, [^6] Nastase and Merlo (2023) show that task-relevant information is more easily accessible in transformer-based sentence embeddings reshaped as two-dimensional arrays, which indicates that patterns are encoded periodically, and are best detected with a 15x15 kernel. followed by a linear layer that compresses the output of the CNN into a latent layer of size 5. The decoder is a mirror-image of the encoder, and unpacks a sampled latent vector into a 32x24 sentence representation.
@@ 4.1.2 Analysis
4.1.3 Electra vs. BERT and RoBERTa, and the price of fine-tuning
We have chosen Electra for the investigations presented here, because with its use of the full context of the sentence to determine whether tokens in the input have been replaced by a generator model, it provides a stronger supervision signal than BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019). To check whether the sentence embeddings produced by the three systems differ in their encoding of chunks, we ran the same experiment on the data encoded with BERT [^7] https://huggingface.co/google-bert/bert-base-multilingual-cased and RoBERTa [^8] https://huggingface.co/FacebookAI/xlm-roberta-base. In terms of F1 score on the task of reconstructing a sentence with the same chunk structure, BERT has a mean over 3 runs of 0.91 (std=0.0346), while RoBERTA has 0.8926 (std=0.0166).
Sentence embeddings are often fine-tuned for specific tasks. We tested two sentence transformer models LaBSE and MPNet [^9] https://huggingface.co/sentence-transformers/LaBSE,https://huggingface.co/sentence-transformers/all-mpnet-base-v2, and obtained an F1 mean of 0.43 (std=0.0336) and 0.669 (std=0.0407) respectively. We chose LaBSE and MPNet because they are two differently tuned models –LaBSE is trained with bilingual sentence pairs with high results on a cross-language sentence retrieval task, MPNet is optimized for sentence similarity – and their representations have the same dimensionality (768) as the transformer models we used. The low results on detecting chunk structure in sentence embeddings after this tuning indicates that in the quest of optimizing the representation of the meaning of a sentence, structural information is lost.
@@ 4.2.2 Analysis
Second, from the results in terms of F1, we note that the two-level process generalizes better from simpler data – learning on type I and type II leads to better results on all test data, with the highest improvement when tested on type III data, which has the highest lexical variation.
Furthermore, the two-level models learned when training on the lexically simpler data perform better when tested on the type III data than the models learned on type III data itself. This result not only indicates that structure information is more easily detectable when lexical variation is less of a factor, but more importantly, that chunk information is separable from other types of information in the sentence embedding, as the patterns detecting it can be applied successfully for data with additional (lexical) variation. [^10] It might appear surprising that the two-level approach leads to lower performance on type III data, particularly when lexical variation had not been an issue for the sentence representation analysis (Section 4.1). The difference comes from the way the instances were formed, on the fly, for the two-level process: the positive sentence to be reconstructed is the same as the input, instead of being a sentence that has the same structure, but different lexical material. This is because all sentences in the sequence have different structures. We think this weakens the (indirect) supervision signal – as the correct answer is distinct from the other options. This is not the case for type I and II data, where, because of the very similar lexical material, the distinction between the correct and incorrect answers reduce to the structure. We plan to confirm this in future work using a pre-trained sentence-level VAE.